OpenStudy (anonymous):

The slope of the line is 2.5. Write the equation of the line in point-slope form, using the coordinates (2,5). Do not use parenthesis on the y side.

OpenStudy (anonymous):

y – y1 = m(x – x1)

OpenStudy (agent0smith):

Point slope form is \[\Large y- y_1 = m(x-x_1)\] m is the slope, 2.5. \[\Large y- y_1 = 2.5(x-x_1)\] Then just plug in your x1 and y1, (2,5)